Find a set of solutions for:
2(x-y)dx+(3x-y-1)dy=0

(2x+y-4)dx+(x-3y+12)dy=0

dy/dx=tanycotx - secycosx
1
Expert's answer
2020-04-04T15:58:47-0400

1)

"2(x-y)dx+(3x-y-1)dy=0\\
y'=\frac{2(y-x)}{3x-y-1}"

Solving the system

"\left \{
\begin{matrix}
 y_0-x_0=0 \\
 3x_0-y_0-1=0 
\end{matrix}\right.\\
\left\{\begin{matrix}
 x_0=\frac{1}{2} \\
 y_0=\frac{1}{2}
\end{matrix}\right."

We shall replace

"\left\{\begin{matrix}
 x=t+\frac{1}{2} \\
 y=z+\frac{1}{2}
\end{matrix}\right.\\
y'=z'"

where "t" is a new independent variable, "z" is a new function.

Substitute into equation

"z'=\frac{2(z+\frac{1}{2}-t-\frac{1}{2})}{3t+\frac{3}{2}-z-\frac{1}{2}-1}\\
z'=\frac{2(z-t)}{3t-z}"

We shall replace

"\frac{z}{t}=u(t), z=t\cdot u, z'= tu'+u"

then

"tu'+u=\frac{2(tu-t)}{3t-tu}\\
tu'+u=\frac{2u-2}{3-u}\\
tu'=\frac{2u-2}{3-u}-u\\
tu'=\frac{2u-2-3u+u^2}{3-u}\\
tu'=\frac{u^2-u-2}{3-u}\\
\frac{3-u}{(u+1)(u-2)}du=\frac{dt}{t}\\
\int(-\frac{4}{3}\frac{1}{u+1}+\frac{1}{3}\frac{1}{u-2})dy=\int\frac{dt}{t}\\
-\frac{4}{3}ln|u+1|+\frac{1}{3}ln|u-2|=ln|t|+\frac{1}{3}ln|c|\\
\frac{u-2}{(u+1)^4}=ct^3, u+1=0"

where "u=\frac{y-\frac{1}{2}}{x-\frac{1}{2}}"

Then solutions of the equation are

"\frac{\frac{y-\frac{1}{2}}{x-\frac{1}{2}}-2}{(\frac{y-\frac{1}{2}}{x-\frac{1}{2}}+1)^4}=c(x-\frac{1}{2})^3,\\
\frac{y-\frac{1}{2}}{x-\frac{1}{2}}+1=0"


2)

"(2x+y-4)dx+(x-3y+12)dy=0\\
y'=-\frac{2x+y-4}{x-3y+12}"

Solving the system

"\left\{\begin{matrix}
 2x_0+y_0-4=0 \\
 x_0-3y_0+12=0
\end{matrix}\right.\\
\left\{\begin{matrix}
 x_0=0 \\
 y_0=4
\end{matrix}\right."

We shall replace

"\left\{\begin{matrix}
 x=t \\
 y=z+4
\end{matrix}\right.\\
y'=z'\\
z'=\frac{2t+z+4-4}{t-3z-12+12}\\
z'=-\frac{2t+z}{t-3z}"

We shall replace

"\frac{z}{t}=u(t), z=t\cdot u, z'= tu'+u"

then

"tu'+u=\frac{2t+tu}{3tu-t}\\
tu'=\frac{2+u}{3u-1}-u\\
tu'=\frac{2+u-3u^2+u}{3u-1}\\
tu'=\frac{-3u^2+2u+2}{3u-1}\\
-\frac{3u-1}{3u^2-2u-2}du=\frac{dt}{t}\\
-\frac{1}{2}\int\frac{d(3u^2-2u-2)}{3u^2-2u-2}du=\frac{dt}{t}\\
-\frac{1}{2}ln|3u^2-2u-2|=ln|t|+\frac{1}{2}ln|c|\\
\frac{1}{3u^2-2u-2}=ct^2\\
3u^2-2u-2=0"

where "u=\frac{y-4}{x}"

Then solutions of the equation are

"\frac{1}{3(\frac{y-4}{x})^2-2\cdot\frac{y-4}{x}-2}=cx^2\\
3(\frac{y-4}{x})^2-2\cdot \frac{y-4}{x}-2=0"


3)

"\frac{dy}{dx}=\tan y\cot x-\sec y\cos x\\
\frac{dy}{dx}=\frac{\sin y}{\cos y}\frac{\cos x}{\sin x}-\frac{1}{\cos y}\cos x\\
\frac{dy}{dx}=\frac{\sin y\cos x-\sin x \cos x}{\sin x\cos y}\\
(\sin x-\sin y)\cos x dx+\sin x \cos y dy=0"

Put "u=\sin y, du=\cos ydy"

Substituting, we get

"(\sin x-u)\cos xdx+\sin x du=0\\
\frac{du}{dx}-u\cdot\frac{\cos x}{\sin x}=-\cos x"

The equation is first - order linear in "u"

The integrating factor is

"I=e^{-\int\frac{\cos x}{\sin x}dx}=e^{-\ln\sin x}=\frac{1}{\sin x}"

Hence

"u\frac{1}{\sin x}=-\int\frac{\cos x}{\sin x}dx=-ln|\sin x|+c"


solve for "u"

"u=-\sin x\ln|\sin x|+c\sin x"

Put "y" back

"\sin y=-\sin x\ln|\sin x|+c\sin x"